OpenStudy (anonymous):

If f(X)=x^2 + 4 and g(y)= 2/√y then (g o f ) (t) =…

OpenStudy (anonymous):

\[g\circ f(t)=g(f(t))=g(t^2+4)=\frac{2}{\sqrt{t^2+4}}\]